Bunker makes two types of briefcase, fabric and leather. The company is currently using a traditional costing system with labor hours as the cost driver but is considering switching to an activity-based costing system. In preparation for the possible switch, Bunker has identified two activity cost pools: materials handling and setup. Pertinent data follow:     

Fabric Case Leather Case
Number of labor hours 15,000 6,000
Number of material moves 672 1,428
Number of setups 32 238


Total estimated overhead costs are $393,300, of which $315,000 is assigned to the materials handling cost pool and $78,300 is assigned to the setup cost pool.

Required:

1.
Calculate the overhead assigned to the fabric case using the traditional costing system based on direct labor hours. (Do not round intermediate calculations and round your final answer to the nearest whole dollar amount.)



2. Calculate the overhead assigned to the fabric case using ABC. (Round activity rates or activity proportions, and intermediate calculations to four decimal places.)



3. Was the fabric case over- or undercosted by the traditional cost system compared to ABC?
Answer #1

1) Traditional costing

Overhead Assigned under traditional costing = 393300/21000*15000 = 280929

2) ABC Costing

Material handling rate = 315000/2100 = 150 per move

Setup cost = 78300/270 = 290 per setup

Overhead assigned to ABC = (672*150+32*290) = 110080

3) Fabric case is over costed under traditional costing system compared to ABC